OpenStudy (anonymous):

Timothy deposited $2,140.30 in a savings account that earns 3.5% simple interest. What will Timothy’s account balance be in 9 months?

OpenStudy (anonymous):

I = PRT I = (2140.30)(.035)(9/12)--9/12 = 3/4 = .75 I = 74.91 (.75) I = 56.18 --this is Timothy's interest for 9 months Timothy's balance will be principle + interest = 2140.30 + 56.18 = 2196.48 I hope I did that correctly....you might need to wait for a second opinion :)

OpenStudy (anonymous):

No its perfect thank you=D your awsome

OpenStudy (anonymous):

your welcome....thats what I am here for :)

OpenStudy (anonymous):

This is from 2 years ago, you guys probably don't even pay attention to this anymore but i have a question, how did you get 74.91 @kelliegirl33